OpenStudy (anonymous):

The lines 2x+y =3 , x+2y = 3 , 2x + y =5, x+ 2y = 5 form the Side of (a) Square (b) Rhombus) (c) Rectangle

OpenStudy (anonymous):

B)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@Lyndsey_Coach_xox hw ?

OpenStudy (anonymous):

Just find the slope here I guess..

OpenStudy (anonymous):

Or distance??

OpenStudy (anonymous):

And don't know finally.. ha ha ha..

OpenStudy (anonymous):

first confirm that they are a set of parallel lines, this can be done by looking at their slope now the difference between (squares and rectangles) and a rhombus is that squares and rectangles have 90 degree angles so in terms of slope, they would be negative reciprocals of each other if they are 90 degree angles, then just check the distances between the intersections of hte line

OpenStudy (anonymous):

I am somewhat right..

OpenStudy (anonymous):

Find slope by easier method using the formula: \[Slope = -\frac{y \; \; coefficient}{x \; \; coefficient}\]

OpenStudy (anonymous):

I think x is above..

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

\[Slope = -\frac{x \; \; coefficient}{y \; \; coefficient}\]

OpenStudy (anonymous):

much easier if you converted into slope intercept form

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

Let @Yahoo! decide..

OpenStudy (anonymous):

And if you are given with two points: \[\frac{y-y_1}{y_2-y_1} = \frac{x-x_1}{x_2-x_1}\]

OpenStudy (anonymous):

Sorry, TWO POINT FORM..

OpenStudy (anonymous):

thxx..Guys this really helps......)

OpenStudy (anonymous):

You are welcome..

OpenStudy (anonymous):

Along with dear..